Bill is 5 years less then 3 times charity's age. The sum of there ages is 79. How old is bill and charity?
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

Bill is 5 years less then 3 times charity's age. The sum of there ages is 79. How old is bill and charity?
----------------------------
Equations::
b = 3c - 5
b + c = 79
----------------------
Substitite for "b" and solve for "c":
3c-5 + c = 79
4c = 84
c = 21 (Charity's age now)
----
Solve for "b":
b + c = 79
b + 21 = 84
b = 63 (Bill's age now)
----------------------
Cheers,
Stan H.
=====================

Answer by richwmiller(17219)

b=3c-5,
b+c=79
b = 58, c = 21